A dog jogs 4m down a hall in 2 seconds and jogs 3m down a perpendicular hall in 1.5 seconds. What was its average speed? What was its average velocity?
1
Expert's answer
2019-08-30T09:34:48-0400

(a)


"{\rm Average\; speed}=\frac{{\rm Total\; distance}}{{\rm Total\; time}}"




"=\frac{4\:\rm{m}+3\:\rm{m}}{2\:\rm{s}+1.5\:\rm{s}}=2\:\rm{m/s}"

(b)


"{\rm Average\; velocity}=\frac{{\rm Total\; displacement}}{{\rm Total\; time}}"

"=\frac{\sqrt{(4\:\rm{m})^2+(3\:\rm{m})^2}}{2\:\rm{s}+1.5\:\rm{s}}=1.4\:\rm{m/s}"